Two airplanes leave an airport at the same time and travel in opposite directions. One plane travels 95 km/h slower than the other. If the two planes are 6556 km apart after 4 hours, what is the rate of each plane?
Answer by jorel1380(3719)

4(x+x-95)=6556
2x-95=1639
2x=1734
x=867
x-95=772..